An income statement for Sams Bookstore for the first quarter of the year is presented below: Sams Bookstore Income Statement For Quarter Ended March 31 Sales Cost of goods sold Gross margin Selling and administrative expenses $1,000,000 665,000 335,000 Selling Administration $107,000 118,000 225,000 Net operating income $ 110,000 On average, a book sells for $50. Variable selling expenses are $4 per book with the remaining selling expenses being fixed. The variable administrative expenses are 3% of sales with the remainder being fixed. The contribution margin for Sams Bookstore for the first quarter is:Which of the following statements concerning direct and indirect costs is NOT true? Multiple Choice A direct cost is one that can be easily traced to the particular cost object. Whether a particular cost is classified as direct or indirect does not depend on the cost object. A particular cost may be direct or indirect, depending on the cost object. The factory managers salary would be classified as an indirect cost of producing one unit of product.During the month of May, direct labor cost totaled $14,680 and direct labor cost was 40% of prime cost. lf total manufacturing costs during May were $74,600, the manufacturing overhead was: Multiple Choice $22,020 $59,920 $36,700 $37,900Mark is an engineer who has designed a telecommunications device. He is convinced that there is a big potential market for the device. Accordingly, he has decided to quit his present job and start a company to manufacture and market the device. Mark purchased a machine two years ago to make experimental boards. The machine will be used to manufacture the new board. The cost of this machine is: Multiple Choice a period cost a differential cost a sunk cost
Answer #1
1)
Sam's Book Store
Unit Sales = $,1000,000/$50          20,000.00 books
Income Statement
Sales $         10,00,000.00
Less: Variable cost
COGS $   6,65,000.00
Variable selling ($4 per book × 20,000 books)          80,000.00
Variable administrative (3% of $1,000,000) $      30,000.00 $           7,75,000.00
Contribution Margin $           2,25,000.00
2)
Option 2 is correct
Costs can be direct and indirect depending on the cost object like product, department, and others such as division, customer, geographic market.
3)
Prime Cost = DM + DL
Manufacturing Cost = DM + DL + MO
DL 14680
Prime Cost = DL /40% = 14680/40% 36700
Manufacturing Cost = Prime Cost + MO
74600 = 36700 + MO
Manufacturing Overhead = 74600 - 36700 37900
4)
Mark is an engineer who has designed a telecommunications device. He is convinced that there is a big potential market for the device. Accordingly, he has decided to quit his present job and start a company to manufacture and market the device.
A sunk Cost
A sunk cost is cost that is incurred in past will not be consider at present.